OpenStudy (mpj4):

How do you interpret this? if x = 2 c + 1, y = -c +3 and z = 2x + 5y - 4, write z in terms of c

OpenStudy (bibby):

plug in the given data and combine t he like terms z is already isolated on the left side of the equation so we don't have to do anything with it

OpenStudy (mpj4):

Can you explain what write z in terms of c more clearly? I didn't quite get that.

OpenStudy (bibby):

what I'd do is: x = 2c + 1 y = -c +3 z = 2x + 5y - 4 z = 2(2c+1) + 5(-c+3) - 4

OpenStudy (mpj4):

Ah, thank you.

OpenStudy (bibby):

distribute the 2 and 5 and combine like terms

OpenStudy (mpj4):

I don't want you to waste any more of your time, the equation was enough. It got the right answer.

OpenStudy (bibby):

it's not a waste of time, I'm just watching a documentary and pissing away time on the internet. this is what I'm here for

OpenStudy (mpj4):

Haha! xD

OpenStudy (bibby):

so yeah. take your time working that stuff out and I'll check your work

OpenStudy (mpj4):

z = -c + 13

OpenStudy (bibby):

\(z = 2(2c+1) + 5(-c+3) - 4\) \(z = 4c+2 + -5c+15 - 4\) \(z = 2 -c+15 - 4\) \(z = -c+13\)

OpenStudy (bibby):

noice

OpenStudy (mpj4):

Alright, I'll be going. Thanks a lot man.

OpenStudy (bibby):

have a good day/afternoon/whatever brah

OpenStudy (mpj4):

Ditto